# CodeCloak: A Method for Evaluating and Mitigating Code Leakage by LLM Code Assistants

### v1

@@ Abstract
LLM-based code assistants are becoming increasingly popular among developers. These tools help developers improve their coding efficiency and reduce errors by providing real-time suggestions based on the developer’s codebase. While beneficial, these tools might inadvertently expose the developer’s proprietary code to the code assistant service provider during the development process. In this work, we propose two complementary methods to mitigate the risk of code leakage when using LLM-based code assistants. The first is a technique for reconstructing a developer’s original codebase from code segments sent to the code assistant service (i.e., prompts) during the development process, enabling assessment and evaluation of the extent of code leakage to third parties (or adversaries). The second is CodeCloak, a novel deep reinforcement learning agent that manipulates the prompts before sending them to the code assistant service. CodeCloak aims to achieve the following two contradictory goals: (i) minimizing code leakage, while (ii) preserving relevant and useful suggestions for the developer. Our evaluation, employing GitHub Copilot, StarCoder, and CodeLlama LLM-based code assistants models, demonstrates the effectiveness of our CodeCloak approach on a diverse set of code repositories of varying sizes, as well as its transferability across different models. In addition, we generate a realistic simulated coding environment to thoroughly analyze code leakage risks and evaluate the effectiveness of our proposed mitigation techniques under practical development scenarios.
Index Terms:
Code assistant, LLM, DRL, Security, Privacy
§§footnotetext: Equal contribution
I Introduction
I-A Machine Learning-Based Code Assistant
The use of large language models (LLMs) for code generation is growing and becoming increasingly useful to the global community [34]. The efficiency of these models has been demonstrated in a number of studies [3]. Code assistance tools, like GitHub Copilot, [^1] https://github.com/features/copilot StarCoder, [^2] https://huggingface.co/bigcode/starcoder and Code Llama [^3] https://ai.meta.com/blog/code-llama-large-language-model-coding/ leverage advanced LLMs to provide real-time suggestions, improving developers’ coding efficiency and reducing errors.
The flow of interaction with a code assistant during the coding process takes place as follows:
  1. 1.
    The developer writes code or describes the desired functionality in code or free text, within an integrated development environment (IDE).
  2. 2.
    The code assistant client, which is often a plugin integrated in the IDE, actively monitors the code being developed. It gathers relevant contextual information, including the developer’s recent interactions in the IDE and segments of code that may be relevant from other code sections and files in order to generate code suggestions.
  3. 3.
    Then, based on the data gathered, the code assistant client generates a prompt that contains relevant information for generating useful suggestions and sends it to the code assistant service (also referred to as model).
  4. 4.
    The code assistant service uses an LLM to generate suggestions about how the code can be completed and sends them back to the code assistant client integrated in the IDE.
  5. 5.
    These suggestions are presented to the developer via the IDE interface. The developer can integrate suitable suggestions in the code.
I-B Privacy and Security Risks
While LLMs have become invaluable in various tasks, their integration and use may inadvertently introduce security risks and expose sensitive data. In the context of code assistants, the use of LLMs may introduce new vulnerabilities to the developed code [22] or unintentionally expose sensitive data (i.e., proprietary code).
Previous studies have addressed three main security and privacy concerns: (1) the risk of inferring whether specific code was used to train an LLM model, known as a membership inference attack [2, 33, 20]; (2) the challenge of preventing proprietary code from being recommended by an LLM-based code assistant, particularly when the model has been trained on similar code, such as code from an open-source repository, by employing code manipulation strategies before the code is released [11]; and (3) the exploration of whether an LLM code assistant can be poisoned during its training/fine-tuning phase, leading to the presentation of intentional suggestions of vulnerable code to developers, thereby introducing vulnerabilities into the developed software [23, 14, 10].
However, while these security and privacy concerns have received media attention, [^4] https://www.forbes.com/sites/siladityaray/2023/05/02/samsung-bans-chatgpt-and-other-chatbots-for-employees-after-sensitive-code-leak/?sh=581383d16078 to the best of our knowledge, the risk and potential extent of private code leakage via prompts transmitted to a third-party code assistant service, along with strategies to mitigate this risk, have not been investigated. We regard this as a significant risk, given that the entire codebase of a proprietary software of an organization could potentially be leaked to an untrusted third-party service provider or an adversary capable of intercepting communication between developers and code assistant services. Therefore, there is a need to thoroughly assess the potential extent of leakage via prompts and devise effective mitigation strategies.
One approach for addressing the privacy concern is to use a local code assistant LLM, which is possible due to the proliferation of open-source models. The use of local models prevents sending prompts outside the network. However, commercial remote alternatives, offer advantages in terms of model freshness, maintenance, performance, ease of use, and integration with well-designed plugins for IDEs. The convenience offered by these commercial services continues to attract developers, despite the privacy concerns associated with remote services.
@@ I-B Privacy and Security Risks
Figure 1: An example of CodeCloak’s performance: On the left, we present an original prompt and the code suggested by StarCoder. On the right, we present the manipulated prompts resulted by applying CodeCloak and the code suggested by StarCoder for the manipulated prompt.
There is a lack of both comprehensive methods for evaluating the extent of such leakage and effective mitigation strategies that can minimize code exposure. In this research we address this gap and propose two complementary methods for mitigating the risk of code leakage.
First, we introduce a technique for reconstructing the developer’s original code from prompts sent from the the IDE to the code assistant service during the development process. This reconstruction method allows code leakage assessment and evaluation when using LLM-based code assistants.
Then, we present CodeCloak, a novel deep reinforcement learning (DRL) agent that manipulates code prompts before they are sent to the code assistant service. Such manipulations may include function removal, renaming variable or function names, or replacing code with textual summarization. CodeCloak aims to achieve two contradictory goals simultaneously: (i) minimizing code leakage and the exposure of the developer’s proprietary code (by introducing manipulations to the prompts), while (ii) preserving relevant and useful suggestions provided by the code assistant.
As an example, in Figure 1 we present an original prompt (left) and a manipulated prompt generated by CodeCloak (right), both provided as an input to the StarCoder code suggestion model. At the bottom, the code suggestion provided by StarCoder for the original prompt and manipulated prompt are presented (marked by the white box). As can be seen, while the manipulated prompt is significantly different from the original one, the response in both cases (suggested code) is similar.
We evaluated our method on various LLM-based code assistants, including CodeLlama and StarCoder, using a unique process simulating real coding environments. The results show that our reconstruction method achieves an average reconstruction success rate of over 90%, indicating the extent of the leakage and demonstrating the need to develop an effective code leakage mitigation mechanism. Our mitigation method, CodeCloak, successfully mitigates the leakage to an average of 44%, while maintaining an average similarity of 71% between the suggestions obtained for the manipulated prompts and the suggestions obtained for the original prompts.
I-C Contributions
The contributions of our work can be summarized as follows:
  • To the best of our knowledge, we are the first to explore the risk of code leakage via the prompts and propose a mitigation.
  • We present a method for reconstructing codebase from code segments; in our usecase, we utilize prompts sent from the IDE to the code assistant model. We utilize the proposed method in order to evaluate the amount and quality of the potential code leakage. However, this method is generic and can be utilized in other usecases, such as in forensics when analyzing code fragment artifacts.
  • We present CodeCloak, a novel method based on a DRL agent, which is aimed at both reducing code leakage and preserving useful suggestions. The DRL agent can be installed as a local agent. CodeCloak is generic and transferable and thus, can be applied to any LLM code assistant. This approach can also be extended to LLMs in other domains.

### v3

# CodeCloak: A Method for Mitigating Code Leakage by LLM Code Assistants

@@ Abstract
LLM-based code assistants are becoming increasingly popular among developers. These tools help developers improve their coding efficiency and reduce errors by providing real-time suggestions based on the developer’s codebase. While beneficial, the use of these tools can inadvertently expose the developer’s proprietary code to the code assistant service provider during the development process. In this work, we propose a method to mitigate the risk of code leakage when using LLM-based code assistants. CodeCloak is a novel deep reinforcement learning agent that manipulates the prompts before sending them to the code assistant service. CodeCloak aims to achieve the following two contradictory goals: (i) minimizing code leakage, while (ii) preserving relevant and useful suggestions for the developer. Our evaluation, employing StarCoder and Code Llama, LLM-based code assistants models, demonstrates CodeCloak’s effectiveness on a diverse set of code repositories of varying sizes, as well as its transferability across different models. We also designed a method for reconstructing the developer’s original codebase from code segments sent to the code assistant service (i.e., prompts) during the development process, to thoroughly analyze code leakage risks and evaluate the effectiveness of CodeCloak under practical development scenarios.
1 Introduction
LLM-Based Code Assistants. The use of large language models (LLMs) for code generation is growing and becoming increasingly useful to the global community Vaithilingam et al. (2022). The efficiency of these models has been demonstrated in a number of studies Chang et al. (2023). Code assistance tools, like Code Llama, [^1] https://ai.meta.com/blog/code-llama-large-language-model-coding/ GitHub Copilot, [^2] https://github.com/features/copilot and StarCoder, [^3] https://huggingface.co/bigcode/starcoder leverage advanced LLMs to provide real-time suggestions, improving developers’ coding efficiency and reducing errors. The flow of interaction with a code assistant during the coding process takes place as follows: First, the developer writes code or describes the desired functionality in code or free text, within an integrated development environment (IDE). The code assistant client, which is often a plugin integrated in the IDE, actively monitors the code being developed. It gathers relevant contextual information, including the developer’s recent interactions in the IDE and segments of code that may be relevant from other code sections and files in order to generate code suggestions. Then, based on the data gathered, the code assistant client generates a prompt that contains relevant information for generating useful suggestions and sends it to the code assistant service (also referred to as the model). The code assistant service uses an LLM to generate suggestions about how the code can be completed and sends them back to the code assistant client. Finally, these suggestions are presented to the developer via the IDE interface. The developer can integrate suitable suggestions in the code.
Privacy and Security Risks. While LLMs have become invaluable in various tasks, their integration and use may inadvertently introduce security risks and expose sensitive data. In the context of code assistants, the use of LLMs may introduce new vulnerabilities to the developed code Pearce et al. (2022) or unintentionally expose sensitive data (i.e., proprietary code). Previous studies have addressed three main security and privacy concerns: (1) the risk of inferring whether specific code was used to train an LLM model, known as a membership inference attack Carlini et al. (2021); Sun et al. (2022); Niu et al. (2023); (2) the challenge of preventing proprietary code from being recommended by an LLM-based code assistant, particularly when the model has been trained on similar code, such as code from an open-source repository, by employing code manipulation strategies before the code is released Ji et al. (2022); and (3) exploration of whether an LLM code assistant can be poisoned during its trainingphase, leading to the presentation of intentional suggestions of vulnerable code to developers, thereby introducing vulnerabilities into the developed software Perry et al. (2022); Khoury et al. (2023); Inan et al. (2021). However, while these security and privacy concerns have received research and media attention [^4] https://www.forbes.com/sites/siladityaray/2023/05/02/samsung-bans-chatgpt-and-other-chatbots-for-employees-after-sensitive-code-leak/?sh=581383d16078, to the best of our knowledge, the risk and potential extent of private code leakage via prompts transmitted to a third-party code assistant service, along with strategies to mitigate this risk, have not been investigated. We regard this as a significant risk given that the entire codebase of an organization’s proprietary software could potentially be leaked to an untrusted third-party service provider or an adversary capable of intercepting communication between developers and code assistant services. One approach for addressing the privacy code leakage concern is to use a local code assistant LLM, which is possible due to the proliferation of open-source models. The use of local models prevents sending prompts outside the network. However, commercial remote alternatives offer advantages in terms of model freshness, maintenance, performance, ease of use, and integration with well-designed plugins for IDEs. The convenience of these commercial services continues to attract developers, despite the privacy concerns associated with remote services.
@@ 1 Introduction
Figure 1: An example of CodeCloak’s performance. On the left, we present an original prompt and the code suggested by StarCoder. On the right, we present the manipulated prompts obtained by applying CodeCloak and the code suggested by StarCoder for the manipulated prompt.
Goal. There is a lack of comprehensive methods that effectively mitigate code exposure. In this research, we address this gap by proposing CodeCloak, a novel deep reinforcement learning (DRL) agent that manipulates code prompts before they are sent to the code assistant service. Such manipulations may include function removal, renaming variable or function names, or replacing code with textual summarization. CodeCloak aims to achieve two contradictory goals simultaneously: (i) minimizing code leakage and exposure of the developer’s proprietary code (by introducing manipulations to the prompts), while (ii) preserving relevant and useful suggestions provided by the code assistant to the developer. As an example, in Figure 1 we present an original prompt (left) and a manipulated prompt generated by CodeCloak (right), both provided as input to the StarCoder code suggestion model. At the bottom, the code suggestions provided by StarCoder for the original prompt and manipulated prompt are presented (outlined in white). As can be seen, while the manipulated prompt is significantly different from the original one, the response in both cases (suggested code) is similar. We also introduce a technique for reconstructing the developer’s original code from prompts sent from the IDE to the code assistant service during the development process. This reconstruction technique enables the assessment and evaluation of code leakage when using LLM-based code assistants. We evaluated our method on various LLM-based code assistants, including Code Llama and StarCoder, using a unique process simulating real coding environments. In our evaluation, our mitigation method, CodeCloak, was shown to successfully reduce code leakage to an average of  40% while maintaining an average of  70% similarity in terms of the CodeBLEU metric, between the suggestions obtained for the manipulated prompts and those for the original prompts.